# Ab initio uncertainty quantification in scattering analysis of microscopy

### v2

@@ VII Discussion
Here, we use DDM [22, 23] as an example to illustrate how building a probabilistic model can connect the original estimator from standard DDM and to improve conventional loss-minimization methods through the maximum marginal likelihood estimator. We show that the estimation in DDM analysis is equivalent to minimizing the temporal variogram of the projected intensity in the reciprocal space based on our probabilistic model. Compared to tracking-based algorithms, DDM eliminates the need to track individual features, but selecting a range of wavevectors is still typically required to minimize the loss function, which can differ in a case-by-case manner. With the probabilistic model of data, we derived the maximum marginal likelihood estimator of the parameters, which optimally weighs information at each wavevector, removing the need for selecting a wavevector range to analyze. By evoking the generalized Schur algorithm for Toeplitz covariance and further reducing data by truncating the high-frequency wavevectors, we can accelerate the computation by around $10^{5}$ times for a microscopy video of regular size, compared to directly computing the likelihood function, allowing almost real-time analysis. We have implemented the AIUQ approaches, and shared them as publicly available software packages available in R and MATLAB [32, 33].
@@
Through a variety of simulated and real experiments of both isotropic and anisotropic processes, we found that the tuning-free AIUQ approach achieves a high estimation accuracy of model parameters and MSDs which were not seen before, justifying the efficiency in integrating the information at different wavevectors by the likelihood function. The $95\%$ confidence intervals of MSDs from the AIUQ estimation are typically narrow yet they cover the true MSDs for most $\Delta t$’s, indicating precise uncertainty quantification. Furthermore, using either the maximum likelihood value or predictive error, our method is able to correctly identify the true model amongst a few possible candidates using imaging data. This aspect had not been previously explored within this context. The key aspect of the AIUQ approach is that it removes the need for selecting wavevector range and weighing information on each wavevector in a case by case manner. This feature is particularly useful to be integrated into high-throughput experiments for automatically determining the gelation time at various experimental conditions.
@@
We outline a few future direction that will overcome the limitation of the AIUQ approaches in the implemented packages. First, we assume that a physical model of the process and equivalently the parametric form of the ISF is known, similar to most conventional DDM analyses in fitting the image structure function. However, this can be restrictive for scenarios where the underlying model governing the dynamic process is unknown. A robust way for estimating the model-free ISF is of great interest, whereas the ISF may contain $n-1$ parameters for a microscopy video with $n$ time frames. Conventional model-free approaches in DDM proceed by directly inverting the ISF at a selected wavevector range separately at each lag time, inevitably truncating lag time range that cannot be reliably analyzed. Consequently, one can be left with a range of $\Delta t$’s much shorter than entire lag time range. Rather, the ISF approximated by MSD can be optimized using the likelihood function in Eq. (9) to efficiently weigh all information at different wavevectors and lag time. To overcome the challenge of optimizing a large number of MSD on all lag times, one may optimize the MSD at a few lag time, then reconstruct the MSD on all lag times by Gaussian process regression [95]. The use of the likelihood function in Eq. (9) can substantially improve the efficiency of a similar approach in [44] by minimizing the loss based on the image structure-function. The model-free version of the AIUQ approach can prove valuable in this context by generating smooth MSD, facilitating the connection of MSD data to frequency-dependent viscoelastic moduli using the GSER [37]. Second, we find that even a small drift can lead to inconsistent estimation for slow dynamics, such as the motion close to the gelation point. It is important to integrate a model containing drift in the ISF. Third, exploring methods for quantifying the uncertainty is an important avenue of future exploration. Merely increasing number of pixels can generate more data but not more information. Instead, increasing the number of florescent probes or extending the time ranges supplies more information to estimate the ISFs. This leads to difficulty in determining the effective sample size in quantifying the uncertainty by asymptotic analysis. Adopting a Bayesian approach that incorporates the number of particles may offer a more dependable means of assessing uncertainty in inverse estimation. Finally, the estimation of a mixture or a distribution of the particles with different sizes can be important for some scenarios, such as aggregation of probes or during polymer degradation processes.
@@ VIII Acknowledgements
This project is partially supported by the National Science Foundation under Grant No. DMS-2053423. MG acknowledges the supplement to National Science Foundation Grant No. 2119663. YH acknowledges partial support from the UC Multicampus Research Programs and Initiatives (MRPI) program under Grant No. M23PL5990 and from the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Grant No. DMR-1933487 (NSF BioPACIFIC MIP). YL acknowledges support from the Biological Sciences Postdoctoral Fellowship Program from Yale Provost’s office. The authors thank Matthew Helgeson, Ryan McGorty, Rae Robertson-Anderson and Megan Valentine for helpful discussions.
@@
$\qquad\mathbb{E}\left[\frac{1}{M}\sum_{j\neq j^{\prime}}\exp\left(i\mathbf{q}\cdot(\mathbf{x}_{j}(t+\Delta t)-\mathbf{x}_{j^{\prime}}(t))\right)\right]$
@@
$=$ $\exp\left(-\frac{q^{2}\Delta x^{2}(\Delta t)}{4}\right),$
where $q^{2}=q^{2}_{1}+q^{2}_{2}$ and the MSD $\Delta x^{2}(\Delta t)=\Delta x^{2}_{1}(t,\Delta t)+\Delta x^{2}_{2}(t,\Delta t)$.
@@
We show another example to indicate the estimation in model-free analysis of DDM depends on the choice of wavevector range [26, 45, 44] using a diffusive process or Brownian motion. Here the MSD $\langle\Delta x^{2}(\Delta t)\rangle$ is obtain by directly inverting the observed image structure function $D(q,\Delta t)$:
$\Delta x^{2}_{est}(q,\Delta t)$ $=\frac{4}{q^{2}}\mbox{ln}\left[\frac{A(q)}{A(q)-D(q,\Delta t)+B(q,\Delta t)}\right].$
We simulate the diffusive process with $500\times 500$ pixels across $500$ frame with diffusion coefficient $\theta=0.5$ in Figure 9. We follow the first approach in fitting $D(q,\Delta t)$ to estimate $A_{j}$ and $\bar{B}$ discussed in Sec III.4. For purely diffusive motion, $\langle\Delta x^{2}(\Delta t)\rangle=4\theta\Delta t$, containing only one parameter with the truth $\theta=0.5$. At $\Delta t=10$, roughly one-third of $q$ are close to the estimate of $\theta$, while at $\Delta t=50$, a much narrower range $q$ produces values close to the true diffusion coefficient. This highlights the need for selecting a wavevector range and weighing the contribution of different $q$’s even for diffusive processes.
@@
We first derive the distribution of $\Delta{\hat{y}}_{re,\mathbf{q}}(t,\Delta t)$ and the distribution of $\Delta{\hat{y}}_{im,\mathbf{q}}(t,\Delta t)$ can be shown similarly. Note $\mathbb{E}[\Delta{\hat{y}}_{re,\mathbf{q}}(t,\Delta t)]=0$. The variance can be computed by
@@
We next derive the probability density of $J$ rings of transformed intensity at the Fourier space in Eq. (9). From Eq. (6), the mean of an $n$-vector $\mathbf{y}_{j^{\prime},re}=(y_{j^{\prime},re}(t_{1}),...,y_{j^{\prime},re}(t_{n}))^{T}$ for any $j^{\prime}\in\mathcal{S}_{j}$ can be computed by the law of total expectation below:
$\mathbb{E}[\mathbf{y}_{j^{\prime},re}]=\mathbb{E}[\mathbb{E}[\mathbf{y}_{j^{\prime},re}\mid\mathbf{z}_{re,j^{\prime}}]]=\mathbb{E}[\mathbf{z}_{re,j^{\prime}}]=\mathbf{0},$
where $\mathbf{z}_{re,j^{\prime}}=(z_{re,j^{\prime}}(t_{1}),...,z_{re,j^{\prime}}(t_{n}))^{T}$ is a vector of zero-mean, real-valued random factors. The covariance of $\mathbf{y}_{re,j^{\prime}}$ can be computed by the law of total covariance:
$\mathbb{V}[\mathbf{y}_{re,j^{\prime}}]$ $=\mathbb{V}[\mathbb{E}[\mathbf{y}_{re,j^{\prime}}\mid\mathbf{z}_{re,j^{\prime}}]]+\mathbb{E}[\mathbb{V}[\mathbf{y}_{re,j^{\prime}}\mid\mathbf{z}_{re,j^{\prime}}]]$
$=\mathbb{V}[\mathbf{z}_{re,j^{\prime}}]+\frac{\bar{B}}{4}\mathbf{I}_{n}=\frac{A_{j}}{4}\mathbf{R}_{j}+\frac{\bar{B}}{4}\mathbf{I}_{n}.$
Note that $\mathbf{y}_{re,j^{\prime}}$ is Gaussian since both $\mathbf{z}_{re,j^{\prime}}$ and $\bm{\epsilon}_{j,re}=(\bm{\epsilon}_{j,re}(t_{1}),...,\bm{\epsilon}_{j,re}(t_{n}))^{T}$ are Gaussian. Furthermore, the random factors $\mathbf{z}_{re,j^{\prime}_{1}}$ and $\mathbf{z}_{re,j^{\prime}_{2}}$ are independent when $j^{\prime}_{1}\neq j^{\prime}_{2}$, and the noise is also independent. Hence the probability density of $\tilde{N}\times n$ matrix $\mathbf{\hat{y}}_{re}=[\mathbf{\hat{y}}_{re}(t_{1}),...,\mathbf{\hat{y}}_{re}(t_{n})]$ follows
@@
$-\frac{S_{j}}{2}\sum_{j^{\prime}\in\mathcal{S}_{j}}\big{(}\mathbf{\hat{y}}_{re,j^{\prime}}^{T}\bm{\Sigma}_{j}^{-1}\mathbf{\hat{y}}_{re,j^{\prime}}+\mathbf{\hat{y}}_{im,j^{\prime}}^{T}\bm{\Sigma}_{j}^{-1}\mathbf{\hat{y}}_{im,j^{\prime}}\big{)},$ (A9)

### v4

@@ VII Discussion
Here, we use DDM [23, 24] as an example to illustrate how building a probabilistic model can connect the original estimator from standard DDM and to improve conventional loss-minimization methods through the maximum marginal likelihood estimator. We show that the estimation in DDM analysis is equivalent to minimizing the temporal variogram of the projected intensity in the reciprocal space based on our probabilistic model. Compared to tracking-based algorithms, DDM eliminates the need to track individual features, but selecting a range of wave vectors is still typically required to minimize the loss function, which can differ in a case-by-case manner. With the probabilistic model of data, we derived the maximum marginal likelihood estimator of the parameters, which optimally weighs information at each wave vector, removing the need for selecting a wave-vector range to analyze. By evoking the generalized Schur algorithm for Toeplitz covariance and further reducing data by truncating the high-frequency wave vectors, we can accelerate the computation by around $10^{5}$ times for a microscopy video of regular size, compared to directly computing the likelihood function, allowing almost real-time analysis. We have implemented the AIUQ approaches, and shared them as publicly available software packages available in R and MATLAB [33, 34].
@@
Through a variety of simulated and real experiments of both isotropic and anisotropic processes, imaged in both fluorescent and bright field modes, we found that the tuning-free AIUQ approach achieves a high estimation accuracy of model parameters and MSDs which were not seen before, justifying the efficiency in integrating the information at different wave vectors by the likelihood function. The $95\%$ confidence intervals of MSDs from the AIUQ estimation are typically narrow yet they cover the true MSDs for most $\Delta t$’s, indicating precise uncertainty quantification. Similar to DDM, our algorithm is robust even in scenarios with small signals on a bright background. Furthermore, using either the maximum likelihood value or predictive error, our method is able to correctly identify the true model amongst a few possible candidates using imaging data. This aspect had not been previously explored within this context. The key aspect of the AIUQ approach is that it removes the need for selecting wave-vector range and weighing information on each wave vector in a case by case manner. This feature is particularly useful to be integrated into high-throughput experiments for automatically determining the gelation time at various experimental conditions.
@@
We outline a few future direction that will overcome the limitation of the AIUQ approaches in the implemented packages. First, we assume that a physical model of the process and equivalently the parametric form of the ISF is known. However, this can be restrictive for scenarios where the underlying model governing the dynamic process is unknown. DDM is also used without fitting [27, 46] by inverting the image structure function separately to estimate the mean square displacement at each $q$, but it still requires the selection of a range of admissible $q$’s. In particular, model-free approaches in DDM proceed by directly inverting the ISF at a selected wave-vector range separately at each lag time [45], inevitably truncating the lag time range that cannot be reliably analyzed. A robust way for estimating the model-free ISF is of great interest. Rather, by adopting the AIUQ framework, the ISF approximated by MSD can be optimized using the likelihood function in Eq. (9) to efficiently weigh all information at different wave vectors and lag time. Using the likelihood function in Eq. (9) can substantially improve the efficiency of a similar approach in [45] by minimizing the loss based on the image structure function, which will facilitate the connection of MSD data to frequency-dependent viscoelastic moduli using the GSER [38]. Second, We find that a small drift is crucial to model for slow dynamics, such as the motion close to the gelation point. The AIUQ framework allows one to include the effect of drifts in modeling ISF, and jointly estimate the intrinsic thermal fluctuation and drifts due to external conditions. Third, the estimation of a mixture or a distribution of the particles with different sizes can be important for some scenarios, such as aggregation of probes or during polymer degradation processes. As the particles with different sizes may have different thermal fluctuations, the correlation patterns modeled in ISF for these particles is also different. It is of interest to include a mixture or a distribution of ISFs in the AIUQ framework for systems with inhomogeneous particle sizes. Lastly, certain particles can undergo rapid photobleaching in some experiments, which can alter the signal-to-noise ratio. Thus, it is of interest to model time-dependent amplitude, $A(\mathbf{q},t)$, and noise parameters $\bar{B}(t)$ for these scenarios.
@@ VIII Acknowledgements
This project is partially supported by the National Science Foundation under Grant No. OAC-2411043. X.L. is supported by the National Science Foundation under Grant No. DMS-2053423. M.G. acknowledges the supplement to National Science Foundation Grant No. 2119663. Y.H. acknowledges partial support from the UC Multicampus Research Programs and Initiatives (MRPI) program under Grant No. M23PL5990 and from the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Grant No. DMR-1933487 (NSF BioPACIFIC MIP). Y.L. acknowledges the support from the National Science Foundation under Grant No. OAC-2411044 and the donors of ACS Petroleum Research Fund under Doctoral New Investigator No. 67734-DNI10. The authors thank Matthew Helgeson, Ryan McGorty, Rae Robertson-Anderson and Megan Valentine for helpful discussions.
@@
$\qquad\mathbb{E}\left[\frac{1}{M}\sum_{m\neq m^{\prime}}\exp\left(i\mathbf{q}\cdot(\mathbf{x}_{m}(t+\Delta t)-\mathbf{x}_{m^{\prime}}(t))\right)\right]$
@@
$=$ $\exp\left(-\frac{q^{2}\langle\Delta x^{2}(\Delta t)\rangle}{4}\right),$
where $q^{2}=q^{2}_{1}+q^{2}_{2}$ and the MSD $\langle\Delta x^{2}(\Delta t)\rangle=\mathbb{E}\left[\Delta x^{2}_{1}(t,\Delta t)+\Delta x^{2}_{2}(t,\Delta t)\right]$.
@@
We show another example to indicate the estimation in model-free analysis of DDM depends on the choice of wave-vector range [27, 46, 45] using a diffusive process or Brownian motion. Here the MSD $\langle\Delta x^{2}(\Delta t)\rangle$ is obtained by directly inverting the observed image structure function $D(q,\Delta t)$:
$\Delta x^{2}_{est}(q,\Delta t)$ $=\frac{4}{q^{2}}\mbox{log}\left[\frac{A(q)}{A(q)-D(q,\Delta t)+B(q,\Delta t)}\right].$
We simulate the diffusive process with $500\times 500$ pixels across $500$ frame with diffusion coefficient $\theta=0.5$ in Figure 9. We follow the first approach in fitting $D(q,\Delta t)$ to estimate $A_{j}$ and $\bar{B}$ discussed in Sec III.4. For purely diffusive motion, $\langle\Delta x^{2}(\Delta t)\rangle=4\theta\Delta t$, containing only one parameter with the truth $\theta=0.5$. At $\Delta t=10$, roughly one-third of $q$ are close to the estimate of $\theta$, while at $\Delta t=50$, a much narrower range $q$ produces values close to the true diffusion coefficient. This highlights the need for selecting a wave-vector range and weighing the contribution of different $q$’s even for diffusive processes.
@@
We first derive the distribution of $\Delta{\hat{y}}_{re,\mathbf{q}}(t,\Delta t)$ for any $\mathbf{q}$, t, and $\Delta t$. The distribution of $\Delta{\hat{y}}_{im,\mathbf{q}}(t,\Delta t)$ can be shown similarly. Note $\mathbb{E}[\Delta{\hat{y}}_{re,\mathbf{q}}(t,\Delta t)]=0$. The variance can be computed by
@@
We next derive the probability density of $J$ rings of transformed intensity at the Fourier space in Eq. (9). From Eq. (6), the mean of an $n$-vector $\hat{\mathbf{y}}_{re,\mathbf{j}^{\prime}}=(\hat{y}_{re,\mathbf{j}^{\prime}}(t_{1}),...,\hat{y}_{re,\mathbf{j}^{\prime}}(t_{n}))^{T}$ for any $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$ can be computed by the law of total expectation below:
$\mathbb{E}[\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}]=\mathbb{E}[\mathbb{E}[\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}\mid\mathbf{z}_{re,\mathbf{j}^{\prime}}]]=\mathbb{E}[\mathbf{z}_{re,\mathbf{j}^{\prime}}]=\mathbf{0},$ (A9)
where $\mathbf{z}_{re,\mathbf{j}^{\prime}}=(z_{re,\mathbf{j}^{\prime}}(t_{1}),...,z_{re,\mathbf{j}^{\prime}}(t_{n}))^{T}$ is a vector of zero-mean, real-valued random factors. The covariance of $\hat{\mathbf{y}}_{re,\mathbf{j}^{\prime}}$ can be computed by the law of total covariance:
$\mathbb{V}[\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}]$ $=\mathbb{V}[\mathbb{E}[\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}\mid\mathbf{z}_{re,\mathbf{j}^{\prime}}]]+\mathbb{E}[\mathbb{V}[\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}\mid\mathbf{z}_{re,\mathbf{j}^{\prime}}]]$
$=\mathbb{V}[\mathbf{z}_{re,\mathbf{j}^{\prime}}]+\frac{\bar{B}}{4}\mathbf{I}_{n}=\frac{A_{j}}{4}\mathbf{R}_{j}+\frac{\bar{B}}{4}\mathbf{I}_{n}.$ (A10)
Note that $\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}$ is Gaussian since both $\mathbf{z}_{re,\mathbf{j}^{\prime}}$ and $\bm{\epsilon}_{j,re}=(\bm{\epsilon}_{j,re}(t_{1}),...,\bm{\epsilon}_{j,re}(t_{n}))^{T}$ are Gaussian. Furthermore, the random factors $\mathbf{z}_{re,j^{\prime}_{1}}$ and $\mathbf{z}_{re,j^{\prime}_{2}}$ are independent when $j^{\prime}_{1}\neq j^{\prime}_{2}$, and the noise is also independent. Hence the probability density of $\tilde{N}\times n$ matrix $\mathbf{\hat{y}}_{re}=[\mathbf{\hat{y}}_{re}(t_{1}),...,\mathbf{\hat{y}}_{re}(t_{n})]$ follows
@@
$-\frac{S_{j}}{2}\sum_{\mathbf{j}^{\prime}\in\mathcal{S}_{j}}\big{(}\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}^{T}\bm{\Sigma}_{j}^{-1}\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}+\mathbf{\hat{y}}_{im,\mathbf{j}^{\prime}}^{T}\bm{\Sigma}_{j}^{-1}\mathbf{\hat{y}}_{im,\mathbf{j}^{\prime}}\big{)},$ (A11)
@@
Lastly, we provide the derivation to prove the unbiasedness of the estimator $A_{est,j}$ in Eq. (10). Note that for any $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$ the mean and covariance of $\mathbf{\hat{y}}_{re,\mathbf{j}^{\prime}}=({\hat{y}}_{re,\mathbf{j}^{\prime}}(t_{1}),...,{\hat{y}}_{re,\mathbf{j}^{\prime}}(t_{n}))^{T}$ and $\mathbf{\hat{y}}_{im,\mathbf{j}^{\prime}}=({\hat{y}}_{im,\mathbf{j}^{\prime}}(t_{1}),...,{\hat{y}}_{im,\mathbf{j}^{\prime}}(t_{n}))^{T}$ are the same. Then, by Eq. (A9) and Eq. (A10), for any $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$ and $t_{k}$ with $1\leq k\leq n$ we have
$\mathbb{E}[|\hat{y}_{\mathbf{j}^{\prime}}(t_{k})|^{2}]$ $=\mathbb{E}[\hat{y}_{re,\mathbf{j}^{\prime}}(t_{k})^{2}+\hat{y}_{im,\mathbf{j}^{\prime}}(t_{k})^{2}]$
$=2\times\left(\frac{A_{j}}{4}+\frac{\bar{B}}{4}\right)$
$=\frac{A_{j}+\bar{B}}{2}.$ (A12)
Then
$\mathbb{E}[A_{est,j}]$ $=\frac{2}{S_{j}n}\sum_{\mathbf{j}^{\prime}\in\mathcal{S}_{j}}\sum^{n}_{k=1}\mathbb{E}[|\hat{y}_{\mathbf{j}^{\prime}}(t_{k})|^{2}]-{\bar{B}}$
$=2\times\left(\frac{A_{j}+\bar{B}}{2}\right)-\bar{B}=A_{j}.$ (A13)
Since the expected value of the estimator $A_{est,j}$ is equal to the underlying true value, the estimator $A_{est,j}$ is unbiased.